OpenStudy (anonymous):

Consider the function g(x)=12 . Then g(7) + g(-7) =_______

OpenStudy (blacksteel):

If g(x) = 12, it means that the result will be 12 for any input x. Thus, g(7) + g(-7) = 12 + 12 = 24